# A Study of Nationality Bias in Names and Perplexity using Off-the-Shelf Affect-related Tweet Classifiers

### v1

@@ Abstract
Valentin Barriere Universidad de Chile – DCC, CENIA, Santiago, Chile vbarriere@dcc.uchile.cl                        Sebastian Cifuentes CENIA Santiago, Chile sebstian.cifuentes@cenia.cl
1 Introduction and Related Work
Figure 1: Overview of the counterfactual example creations. We show examples with sentiment and hate speech for variation of the name "Alexander" and two sentences $S^{1}$ and $S^{n}$. $S^{1}:$I do not like you [PER] you fucking bitch".
Recent trend in Natural Language Processing research, like in works published at conference such as ACL Rogers et al. (2023), is to provide open-source data and models Scao et al. (2022). This practice not only enhances its value for general research purposes but also facilitates the deployment of these models in diverse operational settings by companies or stakeholders. Applications such as customer experience, CV screening, Social Media analyses and moderation are example of applications that will directly impact the users in different ways. For this reason, the models applied at large scale should be scrutinized in order to understand their behavior and should tend to be fair by passing successfully a series of test to reduce their biases toward various target groups.
@@ 1 Introduction and Related Work
Although names are not inherently linked to a specific nationality, research has revealed the presence of nationality biases within them. Delving into this underexplored domain, Venkit et al. (2023) shed light on the influence of demographic on biases associated with countries in language models. An and Rudinger (2023) offer insights into the intricate relationship between demographic attributes and tokenization length, particularly focusing on biases related to first names. Zhu et al. (2023) propose to mitigate name bias by disentangling it from its semantic context in machine reading comprehension tasks. Ladhak et al. (2023) investigate the propagation of name-nationality bias, demonstrating through intrinsic evaluation with templates how names and nationalities are intrinsically linked and how biases manifest as hallucinations. Lastly, Barriere and Cifuentes (2024) showed that using names as proxy works to detect country-related biases depends on the sentence’s language, in multilingual sentiment and stance recognition models.
Contributions
We propose an investigation into biases related to country-specific names in widely used off-the-shelf models (Barbieri et al., 2020, 2022), commonly deployed in production environments for Twitter data. [^1] Regarding the number of monthly download of cardiffnlp models from Barbieri et al. (2020, 2022) in the Huggingface Model Hub at the time of writing. Our analysis reveals distinct biases in sentiment, emotion, and hate speech classifiers, showing a propensity to favor names from certain countries while markedly disfavoring those from less Westernized nations, often by a large margin. Furthermore, we establish a global-level correlation between the perplexity of associated PLMs and model predictions across both known and unknown (i.e., Out-of-Distribution; OOD) languages, demonstrated through examples in English, Basque, and Maori. At a local level, we mitigate the influence of syntax on perplexity by examining the correlation among counterfactual examples generated through minor perturbations. Notably, our findings suggest that the frequency of a name’s occurrence during the training phase directly impacts the sentiment model’s tendency to produce positive outputs, which highly disadvantage the non-English (i.e., OOD) persons in a world where English is widely utilized as pivot language. Our method is unsupervised, moreover it can be applied to any classifier and any dataset.
@@ 2 Method
Table 1: Changes in probability output ($\Delta$) and in percentage of examples in each of the predicted classes.
We first rely on Named Entity Recognition (NER) to create counterfactual examples from the target-domain, specific of target groups. The bias is assessed by quantifying the differences in the model outputs. Second, we ran a series of experiences studying the correlation between the output variations and the perplexity. Figure 1 shows an overview of the bias detection.
@@ 2.1 Perturbation-based Counterfactuals
A set of counterfactual examples are constructed from the target-domain data using a NER system combined with a list of most common names from different countries. Each named entity automatically tagged as person is substituted by a random common name from a specific country. More details are found in the Appendix A.
@@ 2.2 Perplexity and Likelihood
The perplexity of a language model measures the likelihood of data sequences and represents how fluent it is Carlini et al. (2018). In simpler terms, perplexity reflects how unexpected a particular sequence is to the model. A higher perplexity suggests that the model finds the sequence more surprising, while a lower perplexity indicates that the sequence is more likely to occur. We refer to the definition of pseudo-log-likelihood introduced by Salazar et al. (2020), the pseudo-perplexity being the opposite of it. More details are provided in Appendix C. In the following, we will not use the term pseudo- when talking about the pseudo- perplexity or likelihood.
@@ 3.1 Experiments
The second experiment aims to show that the model predictions are in general intricately linked with the perplexity even for unknown languages. We calculate the "global" correlation between perplexity and output probabilities in English and unknown languages such as Maori and Basque, which we obtain using Google Machine Translation. [^2] Our aim is to maintain tweet content while utilizing MT, as they did in Balahur and Turchi (2013). More details in Appendix D.
@@
Experimental Protocol
In order to apply our method to data similar to production data, we collected 8,891 random tweets in English by using the IDs from the Eurotweets dataset (Mozetič et al., 2016). For every sentence $x$, we create $50$ random perturbations of this sentence for each of the target countries. Details can be found in Appendix E.
@@ 3.2 Results
Table 1 provides a comprehensive overview of the impact of country-specific named entities on sentiment, emotion, hate speech, and offensive text classifications across diverse classifiers. Notably, it reveals significant variations in model predictions based on the presence of different country names within textual data. For sentiment analysis, it is striking to observe substantial shifts in sentiment probabilities ($\Delta$) across countries. For instance, countries like India, Turkey or Spain exhibit noteworthy deviations in sentiment probabilities, indicating potential biases in classifier outputs concerning specific national contexts. [^3] This is interesting as Spanish (resp. Indian dialects) are the main foreign languages of migrants in US (resp. UK). The percentages of predicted negative, neutral, and positive sentiments further underscore the nuanced nature of these biases, with certain countries consistently receiving more positive or negative sentiment classifications compared to others. Emotion analysis reveals intriguing patterns in the distribution of predicted emotions across countries. Optimism shows an interesting pattern where the non-English names highly decrease this prediction, up to -33% for Moroccan. It is also notable that Moroccan names provoke a very high increase (60%) of anger predictions at the expense of the other classes. Finally, a similar pattern can be seen for the hate speech and offensive text classifiers. English-speaking countries names highly favor hate speech detection, even as a false positive, compared to other countries. For offensive text detection, there is an increase of 6.1% with counterfactuals using US names and a decrease of 4.9% and 2.1% using Moroccan and Hungarian names.
@@
Table 2 shows the correlations between the perplexity and the labels for Sentiment and Hate speech tasks using tweets from different languages, obtained using Machine Translation. For the hate speech model, the global correlation between the hate speech class and the perplexity is almost close to zero for English data, which is good since showing no spurious pattern between perplexity and hate speech prediction. However, the correlations are higher for the unknown language such as Basque and Maori, where it reaches more than 22%. The model tends to classify as hate speech more easily texts having a higher perplexities, i.e., that are outside the training distribution. For the Sentiment model, the pattern for Basque and Maori language is the same, high positive/negative correlation for the negative/positive class, which means that the less the sentence is similar to the train distribution, the more negative it would be.
@@
Table 3 shows correlations between the relative perplexity of the model and the probabilities of different classes. The results are very different from global correlations. Notably, there is a negative correlation between perplexity and positiveness of the sentiment, which implies that names that are more similar to what was seen during the PLM pre-training will imply a more positive output of the sentiment classifier. This trend is particularly pronounced among English-speaking countries. Due to lack of space, more details and results can be found in Appendix F.
@@ 5 Limitations
First, our method only relies on Named Entities, so it does miss all the implicit hate speech. Nevertheless, it is a system with low recall but high precision as when it detects a change, meaning that the classifier behavior is biased. Second, even if our method slightly perturbates the data from the target distribution, it does not explicitly keep it inside, creating examples that might be a bit outside the distribution of the production data. We think that is the reason why we see a general shift toward a more negative sentiment when comparing perturbated examples and true examples (negative predictions always augment while positive predictions always decrease). It would be more natural to use target-data-specific lexicons, or use a generative model to do the job. However, we think that this is a fair comparison toward all the countries and it can drive a pertinent conclusion on a relative bias between the different countries. Finally, we compare a masked language model, but further experiments are left for future workusing generative models such as flan-T5 (Chung et al., 2022) or Mixtral (Jiang et al., 2024) where the same model computes both label and perplexity, for example using label tokens probabilities to estimate the probabilities (Hegselmann et al., 2023).
@@ Appendix A Counterfactual Examples Creation
We rely on the method proposed by Barriere and Cifuentes (2024) to create counterfactual examples, which were used to detect biases on sentiment (Barbieri et al., 2022) and stance recognition models Barriere and Balahur (2023); Barriere and Jacquet (2022); Barriere et al. (2022).
@@
Country-Specific Entities Gazeeters
Our method is relying on country-specific gazeeters, that can be for different type of named entities: one gazeeter of a specific attribute $A$ from a given country $t$ will contain words related to this country. For example, if the name is the attribute and the country is France, we will obtain the set of the most common French names for man or woman $\mathcal{N}_{\text{France}}=\{\text{Matthieu},\text{Jean},\text{Sophie},...\}$ or last names $\mathcal{L}_{\text{France}}=\{\text{Lepennec},\text{Fourniol},\text{Dubois},...\}$. The proposed method relies on gazeeters that are country-specific, that can be for different type of named entities. For more information, the reader is referred to Appendix B.
Data Perturbation
We use a multilingual off-the-shelf NER system available on the Spacy library AI (2023) and created for social media (named xx_ent_wiki_sm) to identify entities for removal in target-domain data, aligning with the data used during model deployment. These entities, in combination with attributes $\mathcal{A}$, form a dataset for generating contrastive examples $\mathcal{S}^{\prime}=\{S^{\prime}_{1},...,S^{\prime}_{|\mathcal{S}|}\}$ related to specific target groups. The random subtraction process follows Ribeiro et al. (2020) method using simple patterns and the Spacy library AI (2023). We utilized a robust model widely employed in industry. However, given the noisy nature of tweets, it may occasionally miss a name but is more likely to rightfully detect one (with lower recall but higher precision on noisy data). We manually examined 100 examples where a Person (PER) entity was detected in our downloaded data, and found a satisfying precision of the NER to be 88%. Subsequently, our method utilizes as templates examples with detected names (which are pertinent templates if precision is high).
@@
Appendix B Gazeeters
These lexicons were obtained from the Wikidata Query Service. [^5] https://query.wikidata.org/ We also collected the most common names by scrapping Wikipedia pages and added this resource to our code, but we found out it was more straightforward to use the Checklist toolbox from Ribeiro et al. (2020). The authors of Checklist collected, also from Wikidata, common first and last names as well as the associated cities from several countries. This makes a total of 16,771 male first names, 12,737 female first names, 14,797 last names from 194 countries.
Appendix C Pseudo-Likelihood
For a sentence $S=w_{1},w_{2},...,w_{|S|}$, the pseudo-log-likelihood ($PLL$) score given by Eq. 1, can be used for evaluating the preference expressed by an MLM for the sentence $S$.
The Log-Perplexity as defined in Carlini et al. (2018) is the negative log likelihood, hence we use pseudo-log-perplexity as simply the opposite of the PLL. [^6] Contrary to the definition of Salazar et al. (2020) defining it on a complete corpus, summing between all the sentences before passing it to exponential.
@@ Appendix C Pseudo-Likelihood
Appendix D Machine Translation
@@
Appendix E Experimental Protocol
The models used were the ones of Barbieri et al. (2020, 2022) for multilingual sentiment analysis, monolingual hate speech, emotion recognition and offensive text detection: cardiffnlp/twitter-xlm-roberta-base-sentiment, cardiffnlp/twitter-roberta-base-hate, cardiffnlp/twitter-roberta-base-emotion, and cardiffnlp/twitter-roberta-base-offensive. The NER used comes from the Spacy All the English tweets from the EuroTweet dataset were downloaded in June 2020, and a random selection of 20% were used as test set. Experiments were run using Tensorflow 2.4.1 Abadi et al. (2016), transformers 3.5.1 Wolf et al. (2019), a GPU Nvidia RTX-8000 and CUDA 12.0. The 8,891 tweets used in the experiment correspond to a random 10% of the English tweets of the EuroTweets dataset (Mozetič et al., 2016).
Appendix F Local Subjectivity-Perplexity Correlation

### v2

@@ Abstract
Valentin Barriere Universidad de Chile – DCC, CENIA Santiago, Chile vbarriere@dcc.uchile.cl                        Sebastian Cifuentes CENIA Santiago, Chile sebstian.cifuentes@cenia.cl
1 Introduction
Figure 1: Overview of the counterfactual example creations. We show examples with sentiment and hate speech for variation of the name "Alexander" and two sentences $S^{1}$ and $S^{n}$. $S^{1}:$I do not like you [PER] you fucking bitch". The NER is applied to the production data to create templates, which are then filled randomly with most common names from gazeeters of different countries to create a pool of counterfactuals. The discrepancies in probabilities is quantified using metrics such as $\Delta$.
Recent trend in Natural Language Processing research, like in works published at conference such as ACL Rogers et al. (2023), is to provide open-source data and models Scao et al. (2022). This practice not only enhances its value for general research purposes but also facilitates the deployment of these models in diverse operational settings by companies or stakeholders. Applications such as customer experience, CV screening, Social Media analyses and moderation are example of applications that will directly impact the users in different ways. For this reason, the models applied at large scale should be scrutinized in order to understand their behavior and should tend to be fair by passing successfully a series of test to reduce their biases toward various target groups. Past study (Ladhak et al., 2023) showed that PLMs are impacted by names, and Barriere and Cifuentes (2024) proposed a method to quantify this to detect biases of the model toward specific countries, using the country most common names as a proxy. We are showing in this paper that this bias is systematic in several widely-used off-the-shelf classifiers on English data, and propose a method to directly link the bias level with the perplexity of the PLM
Contributions
We propose an investigation into biases related to country-specific names in widely used off-the-shelf models (Barbieri et al., 2020, 2022), commonly deployed in production environments for Twitter data. [^1] Regarding the number of monthly downloads of cardiffnlp models from Barbieri et al. (2020, 2022) in the Huggingface Model Hub at the time of writing ($>$4m for sentiment). Our analysis reveals distinct biases in sentiment, emotion, and hate speech classifiers, showing a propensity to favor names from certain countries while markedly disfavoring those from less Westernized nations, often by a large margin. Furthermore, we establish a global-level correlation between the perplexity of associated PLMs and model predictions across both known and unknown (i.e., Out-of-Distribution; OOD) languages, demonstrated through examples in English, Basque, and Maori. At a local level, we mitigate the influence of syntax on perplexity by examining the correlation among counterfactual examples generated through minor perturbations. Notably, our findings suggest that the frequency of a name’s occurrence during the training phase directly impacts the sentiment model’s tendency to produce positive outputs, which highly disadvantage the non-English (i.e., OOD) persons in a world where English is widely utilized as pivot language. Our method is unsupervised, moreover it can be applied to any classifier and any dataset.
2 Related Work
@@
Although names are not inherently linked to a specific nationality, research has revealed the presence of nationality biases within them. Delving into this underexplored domain, Venkit et al. (2023) shed light on the influence of demographic on biases associated with countries in language models. An and Rudinger (2023) offer insights into the intricate relationship between demographic attributes and tokenization length, particularly focusing on biases related to first names. Zhu et al. (2023) propose to mitigate name bias by disentangling it from its semantic context in machine reading comprehension tasks. Ladhak et al. (2023) investigate the propagation of name-nationality bias, demonstrating through intrinsic evaluation with templates how names and nationalities are intrinsically linked and how biases manifest as hallucinations. Lastly, Barriere and Cifuentes (2024) showed that using names as proxy works to detect country-related biases depends on the sentence’s language, in multilingual sentiment and stance recognition models (Barriere and Balahur, 2023; Barriere and Jacquet, 2022; Barriere et al., 2022).
@@ 3 Method
Table 1: Changes in probability output ($\Delta$) and in percentage of examples in each of the predicted classes, both relative to the original unmodified sentence to compare with the model’s likely real-world production settings.
We first rely on Named Entity Recognition (NER) to create counterfactual examples from the target-domain, specific of target groups, following the methodology of Barriere and Cifuentes (2024). The bias is assessed by quantifying the differences in the model outputs. Second, we ran a series of experiences studying the correlation between the output variations and the perplexity. Figure 1 shows an overview of the bias detection.
@@ 3.1 Perturbation-based Counterfactuals
A set of counterfactual examples are constructed from the target-domain data using a NER system combined with a list of most common names from different countries. Each named entity automatically tagged as person is substituted by a random common name from a specific country. Note that the original entity is conserved, by looking in our gazeeters its corresponding gender. More details are found in the Appendix A.
@@ 3.2 Perplexity and Likelihood
The perplexity of a language model measures the likelihood of data sequences and represents how fluent it is Carlini et al. (2018). In simpler terms, perplexity reflects how unexpected a particular sequence is to the model. A higher perplexity suggests that the model finds the sequence more surprising, while a lower perplexity indicates that the sequence is more likely to occur. We refer to the definition of pseudo-log-likelihood introduced by Salazar et al. (2020), the pseudo-perplexity being the opposite of it. For a sentence $S=w_{1},w_{2},...,w_{|S|}$, the pseudo-log-likelihood ($PLL$) score given by Eq. 1, can be used for evaluating the preference expressed by an MLM for the sentence $S$.
The Log-Perplexity as defined in Carlini et al. (2018) is the negative log likelihood, hence we use pseudo-log-perplexity as simply the opposite of the PLL. [^2] Contrary to the definition of Salazar et al. (2020) defining it on a complete corpus, summing between all the sentences before passing it to exponential. More details are provided in Appendix B. In the following, we will not use the term pseudo- when talking about the pseudo- perplexity or likelihood.
@@ 4.1 Experiments
The second experiment aims to show that the model predictions are in general intricately linked with the perplexity even for unknown languages. We first create datasets in these unknown languages using Machine Translation (MT) in order to preserve the semantic content in-between the different languages, as they did in in Balahur and Turchi (2013). We then calculate the "global" correlation between perplexity and output probabilities in English and unknown languages such as Maori and Basque, which we obtain using Google Translate. [^3] Google MT is based on the LLM PaLM 2 (Google, 2023), which should work reasonably well for these two languages already used in production. More details in Appendix C.
@@
4.2 Experimental Protocol
Gazeeters
We used the dataset collected from Wikidata Query Service. [^4] https://query.wikidata.org/ by the authors of Checklist, composed of common first and last names as well as the associated cities from several countries. This makes a total of 16,771 male first names, 12,737 female first names, 14,797 last names from 194 countries.
NER
We use a multilingual off-the-shelf NER system available on the Spacy library AI (2023) and created for social media (named xx_ent_wiki_sm) to identify entities for removal in target-domain data, aligning with the data used during model deployment.
Perturbation
For every sentence $x$, we create $50$ random perturbations of this sentence for each of the target countries.
Dataset
In order to apply our method to data similar to production data, we collected 8,891 random tweets in English by using the IDs from the Eurotweets dataset (Mozetič et al., 2016). The 8,891 tweets used in the experiment correspond to a random selection of 10% of the English tweets of the EuroTweets dataset (Mozetič et al., 2016) downloaded in June 2020.555No label were used.
Tested Classifiers
The models used were the ones of Barbieri et al. (2020, 2022) for multilingual sentiment analysis, monolingual hate speech, emotion recognition and offensive text detection: cardiffnlp/twitter-xlm-roberta-base-sentiment, cardiffnlp/twitter-roberta-base-hate, cardiffnlp/twitter-roberta-base-emotion, and cardiffnlp/twitter-roberta-base-offensive. Experiments were run using Tensorflow 2.4.1 Abadi et al. (2016), transformers 3.5.1 Wolf et al. (2019), a GPU Nvidia RTX-8000 and CUDA 12.0.
@@ 4.3 Results
[…]
@@
Table 2 shows the correlations between the perplexity and the labels for Sentiment and Hate speech tasks using tweets from different languages, obtained using Machine Translation. For the hate speech model, the global correlation between the hate speech class and the perplexity is almost close to zero for English data, which is good since showing no spurious pattern between perplexity and hate speech prediction. However, the correlations are higher for the unknown language such as Basque and Maori, where it reaches more than 22%. The model tends to classify as hate speech more easily texts having a higher perplexities, i.e., that are outside the training distribution. For the Sentiment model, the pattern for Basque and Maori language is the same, high positive/negative correlation for the negative/positive class, which means that the less the sentence is similar to the train distribution, the more negative it would be. Additional experiments using other languages are confirming the results, and are available in Appendix D.
@@
Table 3 shows correlations between the relative perplexity of the model and the probabilities of different classes. The results are very different from global correlations. Notably, there is a negative correlation between perplexity and positiveness of the sentiment, which implies that names that are more similar to what was seen during the PLM pre-training will imply a more positive output of the sentiment classifier. This trend is particularly pronounced among English-speaking countries. Due to lack of space, more details and results can be found in Appendix E.
@@ 6 Limitations
First, our method only relies on Named Entities, so it does miss all the implicit hate speech. Nevertheless, it is a system with low recall but high precision as when it detects a change, meaning that the classifier behavior is biased. Second, even if our method slightly perturbates the data from the target distribution, it does not explicitly keep it inside, creating examples that might be a bit outside the distribution of the production data. We think that is the reason why we see a general shift toward a more negative sentiment when comparing perturbated examples and true examples (negative predictions always augment while positive predictions always decrease). It would be more natural to use target-data-specific lexicons, or use a generative model to do the job. However, we think that this is a fair comparison toward all the countries and it can drive a pertinent conclusion on a relative bias between the different countries. Another bias induction can also come from the fact that some names can be non gendered in some context, such as Claude as a first-name or Jane as a surname (for a man) that would be tagged as feminine. Co-reference resolution could mitigate this issue, even though we believe it is uncommon. Finally, we compare a masked language model, but further experiments are left for future workusing generative models such as flan-T5 (Chung et al., 2022) or Mixtral (Jiang et al., 2024) where the same model computes both label and perplexity, for example using label tokens probabilities to estimate the probabilities (Hegselmann et al., 2023).
Acknowledgements
The authors thank the reviewers for the various comments that helped to improve the manuscript. This work has been partially funded by National Center for Artificial Intelligence CENIA FB210017, Basal ANID.
@@ Appendix A Counterfactual Examples Creation
Country-Specific Entities Gazeeters
Our method is relying on country-specific gazeeters, that can be for different type of named entities: one gazeeter of a specific attribute $A$ from a given country $t$ will contain words related to this country. For example, if the name is the attribute and the country is France, we will obtain the set of the most common French names for man or woman $\mathcal{N}_{\text{France}}=\{\text{Matthieu},\text{Jean},\text{Sophie},...\}$ or last names $\mathcal{L}_{\text{France}}=\{\text{Lepennec},\text{Fourniol},\text{Denis},...\}$. The proposed method relies on gazeeters that are country-specific, that can be for different type of named entities.
Data Perturbation
The detected entities, in combination with attributes $\mathcal{A}$, form a dataset for generating contrastive examples $\mathcal{S}^{\prime}=\{S^{\prime}_{1},...,S^{\prime}_{|\mathcal{S}|}\}$ related to specific target groups. The random subtraction process follows Ribeiro et al. (2020) method using simple patterns and the Spacy library AI (2023). Even though the model utilized is robust and widely employed in the industry, given the noisy nature of tweets, it may occasionally miss a name but is more likely to rightfully detect one (with lower recall but higher precision on noisy data). We manually examined 100 examples where a Person (PER) entity was detected in our downloaded data, and found a satisfying precision of the NER to be 88%. Subsequently, our method utilizes as templates examples with detected names (which are pertinent templates if precision is high).
Appendix B Pseudo-Likelihood
@@
Appendix C Machine Translation
@@
Appendix D Global Subjectivity-Perplexity Correlation
We extend the experiments of Table 2, using the exact same setting, but with other languges: Dutch, Spanish, Hindi, Malayalam and Turkish. We show the results in Table 4. It is possible to see that the sentiment model is behaving for these "known languages" the same way it behaves with English, with a negative correlations on the negative and positive sentiment and a positive correlation with the neutral sentiment. The behavior that we see for out-of-distribution languages such as Maori or Basque is very different.
Appendix E Local Subjectivity-Perplexity Correlation